To: J. Cook, D. Helms, W. Skinner
cc: Division (M7), FEL Coordination Group
From: F. Dylla
Subject: IR Demo Project Weekly Report, August 31-September 4,
Date: September 4, 1998
This week was devoted to scheduled maintenance/installation activities
and continued preparations for the upcoming high power run. Highlights
for the week include: (1) completion of the refurbishment of the
internal hardware that was to be replaced in the photogun (at
week's, end this hardware was being installed); (2) a successful
check of the alignment of the wiggler viewers and associated hardware;
(3) progress with the installation of the Laser Safety System
hardware in Lab 1.
On Thursday Sept.3, the management of the FEL project gave presentations
to a team of security experts from DOE-ORO and DOE headquarters
concerning security precautions that are in place or planned for
use of the FEL Facility. The visiting team was satisfied with
their review of the project.
Also on Thursday, the Laboratory and the Applied Research Center
hosted a day-long conference on state-wide efforts devoted to
semiconductors. The conference was attended by approximately
100 from industrial, university and government sectors. The program
of the conference was key-noted by the Commonwealth's new Secretary
for Technology, Donald Upson, and also included a presentation
and a tour of the FEL Facility.
A major alignment check and readjustment of the wiggler viewers
was performed to bring them back to the original tolerances set
in April. All of the viewers were found to be within 100 microns
or better horizontally but exhibited a vertical displacement of
typically 600 microns. They were reset to within 100 microns
vertically also. This brings them well within lasing tolerance.
Significant progress on the Laser Safety System for the user labs
occurred. Nearly all hardware is complete for Lab 1 and the software
for the overall system is now complete and in beta testing. It
provides a good one-look status of the safety condition of all
labs in EPICs.
Other tasks that have been completed so far during this down period
Install auxiliary contact for 208 vac to HVPS for Gun for PSS.
Brighten viewer 1F02 by a factor of about 10.
Replace camera for 3F02 viewer.
Install new lens for 3F02 viewer.
Check polarity of the magnets in the 5F region.
Provide power supply for cathode camera.
Replace LSS Master.
Test new fast photodiode.
Modify all ion pumps with 10K resistor to prevent oscillations.
Add power disconnect for VBV1F02 solenoid.
Measure the 4 channel BPM's in the 4F region.
Move PSS status to EPICS.
Move Machine Mode Status chassis to FL01B10.
Add Beam Mode status to EPICS.
Get the signals from both Happek Golay cells into EPICS.
Add totalizing counter to SF6 Recovery System.
Put covers on penetrations that are not below racks.
Install Beam Scraper in 3F region.
Check all racks for good connections to the Ground buss cable.
Add shutoff and drain valves in city water line to vault.
Add 3 additional 240 vac outlets next to work bench.
Add 120 vac quad-outlet box with new circuit next to work bench.
Change 120 vac dual outlet box to a quad next to work bench.
Add 120 vac quad-outlet box with new circuit next to north face of column.
Add 120 vac quad-outlet box next to SF6 tank.
Remove extraneous stuff from stands and girders.
Install repaired 4F BPM.
Align sextupole magnets in 3F and 5F regions.
The photocathode gun was reassembled after carefully repolishing
the ball that supports the cathode, and at this writing (1150
Friday, 4 Sep 98) it was being prepared for pump down. A check
of the previous wafer under microscope revealed some splattered
metal, probably stainless steel, that surely contributed to the
field-emission problems that were constraining gun operations
to well below 350 kV during the final days of the last run in
According to the updated plan for refurbishing the gun, we will
be able to pull beam off the gun on Friday, 18 Sep 98. The basic
plans for the week of 14 Sep are to high-voltage process the gun,
do all possible hot system checks without beam, and (hopefully
that Friday) restore the injector setup with the gun operating
at 350 kV. Commissioning activities would then resume in earnest
on 21 Sep, with the principal focus on high-power lasing with
energy recovery. Activities that do not contribute toward that
focus will be deferred, with the exception of trying to get initial
data on CSR during owl shifts.
Concerning CSR experiments, the plan is to see if there is blatant
emittance growth generated in the first recirculation bend. If
so, parametric studies can begin based on measurements of rms
beam properties in the vicinity of the wiggler. If not, careful
setup of the injector may be necessary in that we will then need
to know with higher confidence the phase space of the beam in
the wiggler region to establish more accurately the initial conditions
for simulations. For example, we would probably need good phase-space
measurements at the injection point of the cryomodule (with Multislit
#2 and Happek #1) as a benchmark of the beam dynamics.
The "kilowatt-related" installation/maintenance task
list has grown longer and was prioritized to reflect the "must
haves" to lase at high cw power with energy recovery. As
is evident from the FEL Installation/Maintenance Activities documented
above, progress in working through the task list continues to
Considerable progress was made to establish a refined Injector
Checkout Procedure and Recirculation/Lasing Procedure. Work on
the latter is continuing based on simulation studies that include
the influence of nonlinearities in the magnetic optics of the
recirculation loop on the energy spread at the energy-recovery